Unless otherwise stated in the problem, assume all bonds pay interest semi-annually 1. Even though bond markets are larger than equity markets, bond markets tend to be less liquid than equity markets. Evaluate the words in italics. True or False? a. True b. False 2. Which of the following statements is most likely FALSE: a. Increasing the number of stocks in a portfolio from 10 to 20 will decrease total risk b. Increasing the number of stocks in a portfolio from 10 to 20 will decrease non-systematic c. Decreasing the number of stocks in a portfolio from 20 to 10 will increase non-systematic risk d. Decreasing the number of stocks in a portfolio from 20 to 10 will increase total risk e. Decreasing the number of stocks in a portfolio from 20 to 10 will increase systematic risk 3. You have $2000 in a savings account today (T-0). Using the Rule of 72, approximate how much longer would it take for your investment to quadruple (i.e. to become worth $8000, or double two times) if you earned 8.0% on your savings vs. 24.0% on your savings at the end of every year? a. 6 years b. 9 years c. 12 years d. 16 years e. l8 years 4. When evaluating a project, IRR is better to use than NPV if when after the initial cash outlay there are both positive and negative future cash flows. Evaluate the underlined words in italics. True or False? Hint: Think non-conventional cash flows. a. True b. False
Answer #1

1. False.

2. Option e is Correct Option. Decreasing number if stocks does not increase systematic risk

3. Quadrupling at 8% = 2* 72/8= 18
Quadrupling at 24% = 2* 72/24= 6

So difference = 18-6 = 12 years

4.False. IRR can give multiple values but NPV always gives a single value.